Two boilers one with superheater and other without superheater are delivering equal 

quantities of steam into a common main. The pressure in the boilers and the main is 16 bar. 

The temperature of the steam from a boiler with a superheater is 350°C and temperature of 

the steam in the main is 200°C. Determine the quality of steam supplied by the other boiler.


1
Expert's answer
2021-06-24T10:32:01-0400

From the steam table, at 16 bars, "h= 1959.7 kJ\/kg; t_s=195.07 ^0C; k_p=2.1 kJ\/kg K"

Let the dryness fraction generated by boiler B be x

Then, moisture in 1 kg of wet steam = "(1-x) kg"

Heat required to evaporate the moisture = "(1-x)L"

Heat required to superheat thedry steam = "k_p(t_2-t_s) kJ\/kg"

Heat gained by wet steam of the boiler B "(1-x)L+k_p(t_2-t_s) kJ\/kg"

Heat lost by superheated steam of boiler A "k_p(t_2-t_1) kJ\/kg"

Heat gained by wet steam of the boiler B= Heat lost by superheated steam of boiler A

"(1-x)L+k_p(t_2-t_s) kJ\/kg =k_p(t_2-t_1) kJ\/kg"

"(1-x)*1959.7+2.1*(235-195.07) kJ\/kg =2.1(350-200) kJ\/kg"

"x=0.88204 \implies 88.204 \%"